A biologist is cultivating a kind of super-bacteria in his laboratory. In a successful experiment, the number of super-bacteria will be doubled. However, the number of super-bacteria will stay the same if the experiment fails.

  1. (a) Suppose now there are exactly 1 million super-bacteria and the experiment has a 50 % chance of failing. Also suppose that the outcomes of individual experiments are independent of each other. What is the probability that there are exactly 16 million super-bacteria after doing 10 experiments? Also give a general expression for the probability that there are exactly b million bacteria after a experiments.

  2. (b) Another biologist is also cultivating such super-bacteria, and he has exactly 1 million super- bacteria to begin with. In his laboratory, the number of super-bacteria will be doubled if the experiment is successful. However, the number of super-bacteria will decrease by a factor of two if the experiment fails. The probability of a successful experiment is 1/2, and that of a failed experiment is also 1/2. Suppose that the outcomes of the individual experiments are independent of each other. What is the probability that there are exactly 4 million bacteria after 10 experiments? Also give a general expression for the probability that there are exactly b million bacteria after a experiments.
Answer #1

(a)

Binomial distribution:

Probability of doubling bacteria in one experiment is, p= 0.5

Here, success is "doubling of bacteria".

For 1 million to become 16 million, the 1 million has to be doubled 4 times.

1x2 =2 (1st time); 2x2 =4 (2nd time); 4x2 =8 (3rd time); 8x2 =16 (4th time). Thus, the number of successes (doubling of bacteria) we require is to happen 4 times in order to have 16 million super-bacteria. (Here, 24 =16 and thus, X =4)

Thus, X=4; n=10 experiments.

Formula: P(X=r) =C(n, r).pr.q(n-r) (where, P =Probability)

P(X=4) =C(10, 4) (0.5)4 (1-0.5)(10-4) =210(0.0625)(0.015625) =0.2051

Thus, the probability that there are exactly 16 million super-bacteria after doing 10 experiments is 0.2051​​​​​​

Now, for the general expression, we can use the above binomial distribution formula where n= number of experiments =a (given); now we need the number of successes, r in terms of b. Observe that in the above solution, we have 24 =16 and so, r =4. But now we have "b" million instead of 16 million. So, 2r =b which implies log2b=r and so, write log2b in place of "r".

Hence, a general expression for the probability that there are exactly "b​​​​​​" million bacteria after "a​​​​​​" experiments is:

P(X= log2b) =C(a, log2b). (p)log2b (1 - p)(a - log2b​​​​​​)